In a hydraulic lift, the maximum gauge pressure is 16.1 atm. If the diameter of the output line is 19.3 cm, what is the heaviest vehicle that can be lifted?
Answer #1

d = 19.3 cm = 0.193 m

pressure p = 16.1 atm = 16.1 * 101325 N/m
Answer #2

area of piston = pi* 0.0965
Answer #3

area of piston,A = pi * r^2 = pi* (19.3/2)
Answer #4

given are

P=16.1 atm

Pressure equals Force / Area

P=F/A

converting 16.1 atm in to N/m

P = 16.1 atm X (101.3 kN/m) / (1atm)

=1630.93kN / m

P=1630930N/m

A = pi * r^2

=3.14*0.0965*0.0965

A=0.0292m^2

P=F/A

F=PA

=1630930*0.0292

=47623.156N

F=Mg

47623.156N=m*9.8

47623.156N/9.8=M

4859.5=M

4.8*10^3kg=M